# by Laurent Busé, Thomas Dedieu

### v2

@@ Abstract
where the coefficients $a_{i}$ are homogeneous in $p^{\prime}$ and $p^{\prime\prime}$ of degrees $d-i$ and $i$ respectively. Therefore, we deduce from Example (2.9) (see also Proposition (2.13)) that $\operatorname{redDisc}(f)$ has degrees $(d-2)(d-3)$ and $(d+2)(d-3)$ in $p^{\prime}$ and $p^{\prime\prime}$ respectively, which concludes the proof. ∎
@@
The degree of $\operatorname{redDisc}(f)$ in $p^{\prime\prime}$ may be computed alternatively as follows. The plane curve $C_{p^{\prime}}:=\mathbf{T}_{p^{\prime}}S\cap S$ has in general a double point at $p^{\prime}$, and what we want is the number of lines in $\mathbf{T}_{p^{\prime}}S$ passing through $p^{\prime}$ and tangent to $C_{p^{\prime}}$ at some other point. This is the number of ramification points of the projection of $C_{p^{\prime}}$ from $p^{\prime}$; the latter is a $(d-2):1$ map $\bar{C}_{p^{\prime}}\to\mathbf{P}^{1}$, where $\bar{C}_{p^{\prime}}$ denotes the normalization of $C_{p^{\prime}}$ at $p^{\prime}$, so it follows from the Riemann–Hurwitz formula that the number of ramification points equals
@@
We were not able to find, on the other hand, a geometric argument to compute the degree in $p^{\prime}$ in a comparable fashion. We wonder wether there is an explanation to why this degree $(d-2)(d-3)$ is so nice, in particular in its role in (4.1.1). It is conceivable that it has something to do with the degree of the dual to a smooth plane curve of degree $d-2$.
@@
Let $p\in\mathbf{P}^{3}$ be a general point. The locus of those points $p^{\prime}\in S$ such that $(p^{\prime},p)\in\mathcal{T}_{S}$ (equivalently, $p\in\mathbf{T}_{p^{\prime}}S$) is the apparent boundary $\mathrm{D}^{p}S\cap S$. Among these points $p^{\prime}$, the locus of those $p^{\prime}$ for which the line $\langle p,p^{\prime}\rangle$ is a bitangent to $S$ is cut out by $V(R)$, where $R$ is the polynomial of Proposition (4.2); it is therefore a complete intersection in $\mathbf{P}^{3}$ of type $(d,d-1,(d-2)(d-3))$. One concludes by observing that there are two points $p^{\prime}$ for every bitangent line to $S$ passing through $p$. ∎
@@
Consider an integer $n>3$. Let $V$ be a smooth hypersurface of degree $d\geqslant n+1$ in $\mathbf{P}^{n}$, and $p$ be a general point of $\mathbf{P}^{n}$. The number of lines passing through $p$ and with two contact points of orders $2$ and $n-1$ respectively is
@@
The family of lines in $\mathbf{P}^{n}$ passing through $p$ has dimension $n-1$: indeed, a line containing $p$ is uniquely determined by the datum of a point in the projective quotient $\mathbf{P}^{n}/p$555 this abusive but suggestive notation has the following meaning: if $\mathbf{P}^{n}=\mathbf{P}(\mathbf{k}^{n+1})$, and $\hat{p}\in\mathbf{k}^{n+1}$ represents $p\in\mathbf{P}^{n}$, then $\mathbf{P}^{n}/p$ stands for the $(n-1)$-dimensional projective space $\mathbf{P}\bigl{(}\mathbf{k}^{n+1}/\langle\hat{p}\rangle\bigr{)}$. On the other hand an order $2$ contact (i.e., an ordinary tangency) is $1$ condition, and an order $n-1$ contact is $n-2$ conditions. It is thus expected that the number of lines as in the theorem be finite.
@@
the pair $(p^{\prime},p^{\prime\prime})$ is in $\mathcal{T}_{V}$ if and only if $\langle p^{\prime},p^{\prime\prime}\rangle$ is a line, with contact of order $n-1$ with $V$ at $p^{\prime}$, see [A, Theorem **]. For $(p^{\prime},p^{\prime\prime})\in\mathcal{T}_{S}$, we consider the homogeneous polynomial in $(\alpha,\beta)$
@@
and denote by $\vphantom{f}{}^{\flat}\kern-1.00006ptf(\alpha,\beta)$ the degree $d-n+1$ polynomial between parentheses in the last expression. The line $\langle p^{\prime},p^{\prime\prime}\rangle$ has an additional tangent point with $V$ if and only if the discriminant of $\vphantom{f}{}^{\flat}\kern-1.00006ptf$ vanishes. For all $i=n-1,\ldots,d$, $\mathrm{D}^{{p^{\prime\prime}}^{i}}F(p^{\prime})$ has degree $i$ in $p^{\prime\prime}$ and $d-i$ in $p^{\prime}$. Therefore, by Example (2.9), $\mathrm{Disc}(\vphantom{f}^{\flat}\kern-1.00006ptf)$ has degree $(d-n+1)(d-n)$ in $p^{\prime}$.
Now, the locus of those points $p^{\prime}\in V$ such that $(p^{\prime},p)\in\mathcal{T}_{V}$, i.e., such that the line $\langle p,p^{\prime}\rangle$ has contact of order $n-1$ with $V$ at $p^{\prime}$, is the intersection $V\cap\mathrm{D}^{p}V\cap\cdots\cap\mathrm{D}^{p^{n-2}}$ (see [A, Theorem **], or recall [A, Display **]), which is the complete intersection of $n-1$ hypersurfaces of degrees $d,d-1,\ldots,d-n+2$ respectively. By the first part of the proof, those $p^{\prime}$ such that the line $\langle p,p^{\prime}\rangle$ has an additional contact of order $2$ are cut out in this complete intersection by an $n$-th equation, of degree $(d-n+1)(d-n)$. One thus gets a 00-dimensional complete intersection of degree
@@
The family of codimension $2$ linear subspaces $\Lambda\subseteq\mathbf{P}^{n}$ passing through $n-2$ points $p_{1},\ldots,p_{n-2}$ in linear general position is isomorphic to $\mathbf{P}^{2}$: indeed, such a linear subspace $\Lambda$ is uniquely determined by the datum of a point in the $2$-dimensional projective quotient
@@
(i.e., $\mathbf{P}\bigl{(}\mathbf{k}^{n+1}/\langle\hat{p}_{1},\ldots,\hat{p}_{n-2}\rangle$, see footnote 5). Besides, the Grassmannian of codimension $2$ linear subspaces in $\mathbf{P}^{n}$ has dimension $2(n-1)$, and in this codimension passing through a given point is given by two independent linear conditions.
@@
The degree of the curve $C\subseteq\mathbf{P}^{2}\cong\mathbf{P}^{n}/\langle p_{1},\ldots,p_{n-2}\rangle$ of simply tangent codimension $2$ subspaces is the number of its intersection points with a given line in $\mathbf{P}^{2}$; it is also the number of $(n-2)$-planes $\Lambda\subseteq\mathbf{P}^{n}$ tangent to $V$, passing through $p_{1},\ldots,p_{n-2}$, and contained in a fixed hyperplane $H$ which itself passes through $p_{1},\ldots,p_{n-2}$.
@@
The hypersurface in $\mathbf{P}^{n}$ defined by this homogeneous polynomial is the cone of vertex $\langle p_{1},\ldots,p_{n-2}\rangle$ swept out by those $(n-2)$-planes $\langle p,p_{1},\ldots,p_{n-2}\rangle$ that are tangent to $V$. It projects from $\langle p_{1},\ldots,p_{n-2}\rangle$ to the curve $C$ in $\mathbf{P}^{n}/\langle p_{1},\ldots,p_{n-2}\rangle\cong\mathbf{P}^{2}$, hence $C$ has degree $d(d-1)^{n-2}$.
As we will see in the proof of Theorem (4.1), the tangency points with $V$ of the $(n-2)$-planes parametrized by $C$ form the complete intersection curve $V\cap\mathrm{D}^{p_{1}}V\cap\cdots\cap\mathrm{D}^{p_{n-2}}V$, which has degree $d(d-1)^{n-2}$ and projects birationally to $C$ from $\langle p_{1},\ldots,p_{n-2}\rangle$; this gives another way of computing the degree of $C$.
@@
We continue with the setup introduced in (4.2) above, but now we assume that $p$ sits on $V$ and is such that the $(n-2)$-plane $\langle p,p_{1},\ldots,p_{n-2}\rangle$ is tangent to $V$ at $p$ itself; equivalently, $p$ sits on the complete intersection curve $V\cap\mathrm{D}^{p_{1}}V\cap\cdots\cap\mathrm{D}^{p_{n-2}}V$666this equivalence needs the smoothness of $V$; the fact that this is a complete intersection is equivalent to the fact that tangent subspaces through $p_{1},\ldots,p_{n-2}$ form a curve. Under this assumption, the polynomial $f$ takes the form
@@
The $(n-2)$-plane $\langle p,p_{1},\ldots,p_{n-2}\rangle$ is bitangent to $V$ if and only if the reduced discriminant $\operatorname{redDisc}_{d}^{2}(f)$ vanishes. Thus, the locus of points $p\in V$ such that $\langle p,p_{1},\ldots,p_{n-2}\rangle$ is tangent to $V$ in two points including $p$ is the complete intersection of the curve $V\cap\mathrm{D}^{p_{1}}V\cap\cdots\cap\mathrm{D}^{p_{n-2}}V$ with the hypersurface defined by $\operatorname{redDisc}_{d}^{2}(f)$ as a polynomial in $p$777again, the fact that this is indeed a complete intersection is equivalent to the fact that bitangent subspaces through $p_{1},\ldots,p_{n-2}$ are finitely many, and each of those is tangent to $V$ at only finitely many points; multiple points in this complete intersection will occur if some bitangent subspace has more than two ordinary tangency points with $V$. The latter is homogeneous of degree $d(d-1)^{n-2}-(3n-5)d+6(n-2)$ by Proposition (2.13.1). Therefore, the locus of tangency points of bitangent $(n-2)$-planes through $p_{1},\ldots,p_{n-2}$ has degree
@@
The result follows, since each bitangent $(n-2)$-plane is tangent in two points. ∎
@@
We shall see that those points $p\in S$ such that there is a line intersecting $S$ with multiplicity strictly greater than $3$ at $p$ is a curve $Fl(S)$, cut out on $S$ by a polynomial of degree $2411d-$ 24. We call this curve (respectively polynomial) the flecnodal curve (respectively polynomial) of $S$.
@@
The following statement is definitely a result in reduced elimination theory, although it does not strictly fit in the framework of Section 1.
@@
The locus of those $q\in\mathbf{P}^{3}$ such that $\mathrm{V}(F_{q},G_{q},H_{q})$ contains a point in addition to $q$ counted with multiplicity $\lambda\lambda^{\prime}\lambda^{\prime\prime}$ is the zero locus of a homogeneous polynomial of degree
@@
Computing degrees, one sees that $R$ is homogeneous of degree 00 in the coefficients of $L$, i.e., it does not depend on $L$. It follows that $V(F_{q},G_{q},H_{q})$ has positive dimension if and only if $R(F_{q},G_{q},H_{q})=0$. Eventually, one computes the degrees of $R(F_{q},G_{q},H_{q})$ using the identity (5.2.1). ∎
@@
Let $S$ be a smooth surface in $\mathbf{P}^{3}$ of degree $d$. There exists a homogeneous polynomial $Fl$ of degree $2411d-$ 24, such that the locus of points $p\in S$ such that there is a line intersecting $S$ with multiplicity at least $4$ in $p$ is cut out on $S$ by $\mathrm{V}(Fl)$.
@@
Let $p\in S$. It follows from [A, Thm. **] that there is a line intersecting $S$ with multiplicity at least $4$ in $p$ if and only if the three polar hypersurfaces $\mathrm{D}_{p}S,\mathrm{D}_{p^{2}}S,\mathrm{D}_{p^{3}}S$ (respectively, the tangent plane, the polar quadric, and the polar cubic of $S$ at $p$) have a common point besides $p$, and that this in turn is equivalent to their having a whole line in common.
On the other hand, $\mathrm{D}_{p}S,\mathrm{D}_{p^{2}}S,\mathrm{D}_{p^{3}}S$, which have degrees $1,2,3$ respectively, intersect with multiplicity $6$ at $p$ by [A, Cor. **]. We are therefore in a position to apply Proposition (5.2), with $(\lambda,\mu)$, $(\lambda^{\prime},\mu^{\prime})$, $(\lambda^{\prime\prime},\mu^{\prime\prime})$ equal to $(1,d-1)$, $(2,d-2)$, $(3,d-3)$ respectively. The result follows. ∎

### v3

@@
where the coefficients $a_{i}$ are homogeneous in $p^{\prime}$ and $p^{\prime\prime}$ of degrees did-i and $i$ respectively. Therefore, we deduce from Example (2.10) (see also Proposition (2.14)) that $\operatorname{redDisc}(f)$ has degrees (d2)(d3)(d-2)(d-3) and (d+2)(d3)(d+2)(d-3) in $p^{\prime}$ and $p^{\prime\prime}$ respectively, which concludes the proof. ∎
@@
The degree of $\operatorname{redDisc}(f)$ in $p^{\prime\prime}$ may be computed alternatively as follows. The plane curve $C_{p^{\prime}}:=\mathbf{T}_{p^{\prime}}S\cap S$ has in general a double point at $p^{\prime}$, and what we want is the number of lines in $\mathbf{T}_{p^{\prime}}S$ passing through $p^{\prime}$ and tangent to $C_{p^{\prime}}$ at some other point. This is the number of ramification points of the projection of $C_{p^{\prime}}$ from $p^{\prime}$; the latter is a (d2):1(d-2):1 map $\bar{C}_{p^{\prime}}\to\mathbf{P}^{1}$, where $\bar{C}_{p^{\prime}}$ denotes the normalization of $C_{p^{\prime}}$ at $p^{\prime}$, so it follows from the Riemann–Hurwitz formula that the number of ramification points equals
@@
We were not able to find, on the other hand, a geometric argument to compute the degree in $p^{\prime}$ in a comparable fashion. We wonder wether there is an explanation to why this degree (d2)(d3)(d-2)(d-3) is so nice, in particular in its role in (4.1.1). It is conceivable that it has something to do with the degree of the dual to a smooth plane curve of degree d2d-2.
@@
Let $p\in\mathbf{P}^{3}$ be a general point. The locus of those points $p^{\prime}\in S$ such that $(p^{\prime},p)\in\mathcal{T}_{S}$ (equivalently, $p\in\mathbf{T}_{p^{\prime}}S$) is the apparent boundary $\mathrm{D}^{p}S\cap S$. Among these points $p^{\prime}$, the locus of those $p^{\prime}$ for which the line $\langle p,p^{\prime}\rangle$ is a bitangent to $S$ is cut out by $V(R^{p})$, where $R^{p}=R(\_,p)$ with $R$ the polynomial of Proposition (4.2); it is therefore a complete intersection in $\mathbf{P}^{3}$ of type (d,d1,(d2)(d3))(d,d-1,(d-2)(d-3)). One concludes by observing that there are two points $p^{\prime}$ for every bitangent line to $S$ passing through $p$. ∎
@@
Consider an integer $n>3$. Let $V$ be a smooth hypersurface of degree $d\geqslant n+1$ in $\mathbf{P}^{n}$, and $p$ be a general point of $\mathbf{P}^{n}$. The number of lines passing through $p$ and with two contact points with $V$ of respective orders $2$ and n1n-1 is
@@
[…]
@@
the pair $(p^{\prime},p^{\prime\prime})$ is in $\mathcal{T}_{V}$ if and only if $\langle p^{\prime},p^{\prime\prime}\rangle$ is a line, with contact of order n1n-1 with $V$ at $p^{\prime}$, see [II, Theorem **]. For $(p^{\prime},p^{\prime\prime})\in\mathcal{T}_{S}$, we consider the homogeneous polynomial in $(\alpha,\beta)$
@@
and denote by $\vphantom{f}{}^{\flat}\kern-1.00006ptf(\alpha,\beta)$ the degree dn+1d-n+1 polynomial between parentheses in the last expression. The line $\langle p^{\prime},p^{\prime\prime}\rangle$ has an additional tangent point with $V$ if and only if the discriminant of $\vphantom{f}{}^{\flat}\kern-1.00006ptf$ vanishes. For all $i=n-1,\ldots,d$, $\mathrm{D}^{{p^{\prime\prime}}^{i}}F(p^{\prime})$ has degree $i$ in $p^{\prime\prime}$ and did-i in $p^{\prime}$. Therefore, by Example (2.10), $\mathrm{Disc}(\vphantom{f}^{\flat}\kern-1.00006ptf)$ has degree (dn+1)(dn)(d-n+1)(d-n) in $p^{\prime}$.
Now, the locus of those points $p^{\prime}\in V$ such that $(p^{\prime},p)\in\mathcal{T}_{V}$, i.e., such that the line $\langle p,p^{\prime}\rangle$ has contact of order n1n-1 with $V$ at $p^{\prime}$, is the intersection $V\cap\mathrm{D}^{p}V\cap\cdots\cap\mathrm{D}^{p^{n-2}}$ (see [II, Theorem **]), which is the complete intersection of n1n-1 hypersurfaces of degrees $d,d-1,\ldots,d-n+2$ respectively. By the first part of the proof, those $p^{\prime}$ such that the line $\langle p,p^{\prime}\rangle$ has an additional contact of order $2$ are cut out in this complete intersection by an $n$-th equation, of degree (dn+1)(dn)(d-n+1)(d-n). One thus gets a 0-dimensional complete intersection of degree
@@
The family of codimension $2$ linear subspaces $\Lambda\subseteq\mathbf{P}^{n}$ passing through n2n-2 points $p_{1},\ldots,p_{n-2}$ in linear general position is isomorphic to $\mathbf{P}^{2}$: indeed, such a linear subspace $\Lambda$ is uniquely determined by the datum of a point in the $2$-dimensional projective quotient
@@
(i.e., $\mathbf{P}\bigl(\mathbf{k}^{n+1}/\langle\hat{p}_{1},\ldots,\hat{p}_{n-2}\rangle$, see footnote 7). Besides, the Grassmannian of codimension $2$ linear subspaces in $\mathbf{P}^{n}$ has dimension 2(n1)2(n-1), and in this codimension passing through a given point is given by two independent linear conditions.
@@
The degree of the curve $C\subseteq\mathbf{P}^{2}\cong\mathbf{P}^{n}/\langle p_{1},\ldots,p_{n-2}\rangle$ of simply tangent codimension $2$ subspaces is the number of its intersection points with a given line in $\mathbf{P}^{2}$; it is also the number of (n2)(n-2)-planes $\Lambda\subseteq\mathbf{P}^{n}$ tangent to $V$, passing through $p_{1},\ldots,p_{n-2}$, and contained in a fixed hyperplane $H$ which itself passes through $p_{1},\ldots,p_{n-2}$.
@@
The hypersurface in $\mathbf{P}^{n}$ defined by this homogeneous polynomial is the cone of vertex $\langle p_{1},\ldots,p_{n-2}\rangle$ swept out by those (n2)(n-2)-planes $\langle p,p_{1},\ldots,p_{n-2}\rangle$ that are tangent to $V$. It projects from $\langle p_{1},\ldots,p_{n-2}\rangle$ to the curve $C$ in $\mathbf{P}^{n}/\langle p_{1},\ldots,p_{n-2}\rangle\cong\mathbf{P}^{2}$, hence $C$ has degree $d(d-1)^{n-2}$.
As we will see in the proof of Theorem (4.1), the tangency points with $V$ of the (n2)(n-2)-planes parametrized by $C$ form the complete intersection curve $V\cap\mathrm{D}^{p_{1}}V\cap\cdots\cap\mathrm{D}^{p_{n-2}}V$, which has degree $d(d-1)^{n-2}$ and projects birationally to $C$ from $\langle p_{1},\ldots,p_{n-2}\rangle$; this gives another way of computing the degree of $C$.
@@
We continue with the setup introduced in (4.2) above, but now we assume that $p$ sits on $V$ and is such that the (n2)(n-2)-plane $\langle p,p_{1},\ldots,p_{n-2}\rangle$ is tangent to $V$ at $p$ itself; equivalently, $p$ sits on the complete intersection curve $V\cap\mathrm{D}^{p_{1}}V\cap\cdots\cap\mathrm{D}^{p_{n-2}}V$.888this equivalence needs the smoothness of $V$; the fact that this is a complete intersection is equivalent to the fact that tangent subspaces through $p_{1},\ldots,p_{n-2}$ form a curve. Under this assumption, the polynomial $f$ takes the form
@@
The (n2)(n-2)-plane $\langle p,p_{1},\ldots,p_{n-2}\rangle$ is bitangent to $V$ if and only if the reduced discriminant $\operatorname{redDisc}_{d}^{2}(f)$ vanishes. Thus, the locus of points $p\in V$ such that $\langle p,p_{1},\ldots,p_{n-2}\rangle$ is tangent to $V$ in two points including $p$ is the complete intersection of the curve $V\cap\mathrm{D}^{p_{1}}V\cap\cdots\cap\mathrm{D}^{p_{n-2}}V$ with the hypersurface defined by $\operatorname{redDisc}_{d}^{2}(f)$ as a polynomial in pp.999again, the fact that this is indeed a complete intersection is equivalent to the fact that bitangent subspaces through $p_{1},\ldots,p_{n-2}$ are finitely many, and each of those is tangent to $V$ at only finitely many points; multiple points in this complete intersection will occur if some bitangent subspace has more than two ordinary tangency points with $V$. The latter is homogeneous of degree $d(d-1)^{n-2}-(3n-5)d+6(n-2)$ by (2.14.1). Therefore, the locus of tangency points of bitangent (n2)(n-2)-planes through $p_{1},\ldots,p_{n-2}$ has degree
@@
The result follows, since each bitangent (n2)(n-2)-plane is tangent in two points. ∎
@@
We shall see that those points $p\in S$ such that there is a line intersecting $S$ with multiplicity strictly greater than $3$ at $p$ is a curve $Fl(S)$, cut out on $S$ by a polynomial of degree 11d2411d-24. We call this curve (respectively polynomial) the flecnodal curve (respectively polynomial) of $S$.
@@
The following statement is in the spirit of reduced elimination theory, with the polynomial $R$ (in the notation of the proposition) playing the role of the reduced resultant. Here we are not prescribing individual multiplicities at the point $q$ for the polynomials $F,G,H$ as in the reduced resultant situation, but rather a multiplicity for the ideal they generate; then $R$ vanishes when the subscheme defined by this ideal is larger than what is imposed by the hypothesis, as does $\operatorname{redRes}$ in the reduced resultant situation.
@@
The locus of those $q\in\mathbf{P}^{3}$ such that $\mathrm{V}(F_{q},G_{q},H_{q})$ contains a point in addition to $q$ counted with multiplicity $\lambda\lambda^{\prime}\lambda^{\prime\prime}$ is the zero locus of a homogeneous polynomial $R$ of degree
@@
Computing degrees, one sees that $R$ is homogeneous of degree 0 in the coefficients of $L$, i.e., it does not depend on $L$. It follows that $V(F_{q},G_{q},H_{q})$ has positive dimension if and only if $R(F_{q},G_{q},H_{q})=0$. Eventually, one computes the degrees of $R(F_{q},G_{q},H_{q})$ using the identity (5.2.1). ∎
@@
Let $S$ be a smooth surface in $\mathbf{P}^{3}$ of degree $d$. There exists a homogeneous polynomial $Fl$ of degree 11d2411d-24, such that the locus of points $p\in S$ such that there is a line intersecting $S$ with multiplicity at least $4$ in $p$ is cut out on $S$ by $\mathrm{V}(Fl)$.
@@
Let $p\in S$. It follows from [II, Thm. **] that there is a line intersecting $S$ with multiplicity at least $4$ in $p$ if and only if the three polar hypersurfaces $\mathrm{D}_{p}S,\mathrm{D}_{p^{2}}S,\mathrm{D}_{p^{3}}S$ (respectively, the tangent plane, the polar quadric, and the polar cubic of $S$ at $p$) have a common point besides $p$, and that this in turn is equivalent to their having a whole line in common.
On the other hand, $\mathrm{D}_{p}S,\mathrm{D}_{p^{2}}S,\mathrm{D}_{p^{3}}S$, which have degrees $1,2,3$ respectively, intersect with multiplicity $6$ at $p$ by [II, Cor. **]. We are therefore in a position to apply Proposition (5.2), with $(\lambda,\mu)$, $(\lambda^{\prime},\mu^{\prime})$, $(\lambda^{\prime\prime},\mu^{\prime\prime})$ equal to (1,d1)(1,d-1), (2,d2)(2,d-2), (3,d3)(3,d-3) respectively. The result follows. ∎